Cranberry Inc. purchased a piece of equipment on April 1, 2016 for $180,000. The equipment has a residual value of 30,000 and an estimated useful life of 5 years. Cranberry uses the sum-of-the-years digits to record depreciation expense on this equipment. (a) How much is the depreciation expense in 2018? (11 pts) (b) What is the balance in the Cranberry’s Accumulated Depreciation account for the equipment after recording depreciation expense in 2018? (13 pts) (c) What is the carrying value that Cranberry will report on the Balance Sheet for the equipment in 2018?
Answer #1

a) Depreciation expense in 2018 = 150000*4/15*3/12+150000*3/15*9/12 = 32500

b) Accumulated depreciation = (150000*5/15+150000*4/15+150000*3/15*9/12) = 112500

c) Carrying value in Balance sheet = 150000-112500 = 37500